# Semi-Supervised Learning for Deep Causal Generative Models

### v1

@@ Abstract
Developing models that can answer questions of the form “How would $x$ change if $y$ had been $z$?” is fundamental for advancing medical image analysis. Training causal generative models that address such counterfactual questions, though, currently requires that all relevant variables have been observed and that corresponding labels are available in training data. However, clinical data may not have complete records for all patients and state of the art causal generative models are unable to take full advantage of this. We thus develop, for the first time, a semi-supervised deep causal generative model that exploits the causal relationships between variables to maximise the use of all available data. We explore this in the setting where each sample is either fully labelled or fully unlabelled, as well as the more clinically realistic case of having different labels missing for each sample. We leverage techniques from causal inference to infer missing values and subsequently generate realistic counterfactuals, even for samples with incomplete labels. Code will be available at: https://github.com/yi249/ssl-causal
@@ 1 Introduction
The deployment of deep learning models to real-world applications faces a variety of challenges [16], with many arguing that this is due to lack of causal considerations [4, 18]. A growing research area is the generation of counterfactuals (CFs), the manifestation of a sample in an alternative world where an upstream variable has been changed [17, 21, 10]. Such techniques are particularly useful in medical image analysis, where models are often hampered by lack of diversity in training data [5], so methods to generate realistic synthetic samples from underrepresented classes are critical [13]. Incorporating structural causal equations into a deep learning framework has been shown to provide a powerful tool for counterfactual generation [17]. These ideas were extended by developing a hierarchical VAE structure for greater image fidelity [7]. This method consists, however, of a separately trained generative model and structural causal model, represented by a directed acyclic graph (DAG), and hence, the two components are unable to leverage information from one another during training. Moreover, these methods rely on fully labelled samples, so are unable to use additional data where true values are unavailable for some (or all) variables of the causal graph.
@@ 2 Background
$\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$
where, $\mathbf{V}=\{v_{1},\dots,v_{n}\}$ is the set of endogenous variables of interest, $\mathbf{U}=\{u_{1},\dots,u_{n}\}$ is the set of exogenous (noise) variables, $P(\mathbf{U})$ is the prior distribution over them, and =$\mathcal{F}=\{f_{1},\dots,f_{n}\}$ is a set of functions. Moreover, we assume that each endogenous variable, $x_{i}$, is assigned deterministically by its direct causes, i.e. its parents $\text{pa}_{i}\subseteq\mathbf{V}\setminus\{v_{i}\}$, and the corresponding noise variable, $u_{i}$ via the structural assignments,
@@
$P(\mathbf{V}|\text{do}(v_{i}=a))=\prod_{j\neq i}p(v_{j}|pa_{j},v_{i}=a).$
@@
    Abduction: Use the data to update the prior probability on the exogenous noise $p(\mathbf{U})$ to obtain $p(\mathbf{U}|V)$
@@ 3 Methodology
Overview of our method is shown in Fig. 1. Herein, the endogenous variables consist of image, $x$, and variables, $y$; denoted by $y^{*}$ when observed and by $y^{\prime}$ when predicted. Latent variables $z=z_{1:L}$ make up part of the exogenous noise for $x$, modelled using a hierarchical latent structure inspired by [20]. Our model extends their structure with a predictive part that infers the endogenous variables, $y$, enabling counterfactual generation in the case of missing labels. For clarity, we limit derivations to the case with a single cause variable, $y_{C}$, and effect variable, $y_{E}$, but this can be extended to any finite number of variables with any causal structure. The ELBO loss for fully labelled samples drawn from $\mathcal{D}_{L}$ is $\mathcal{S}(x,y)$:
@@
where $\mathcal{L}(x,y)$ is the ELBO for a conditional VAE and $p_{\phi}(\cdot)$ are (Gaussian) priors. For the unlabelled samples, drawn from $\mathcal{D}_{U}$, we minimise the loss:
@@
Under this construction, the parent predictors, $q_{\phi}(y_{E}|x)$ and $q_{\phi}(y_{C}|x,y_{E})$, are only trained when the parent variables are unobserved. To ensure the model is able to learn a good representation of $y$, we include additional classification terms in the supervised loss [14], giving the total loss to train our model:
@@ 3.0.2 Counterfactual Generation
Once the generative model is trained, we use the predictive component $p(y_{i}|y_{<i})$ to generate counterfactuals, $\tilde{x}$. For abduction we require the structural assignments (1) to be invertible in $u$, so we encode $p(y_{i}|y_{<i})$ as an invertible $g_{y_{<i}}(u),\;u\sim p(u)$, parameterised by the causal parents $y_{<i}$ [17]. Counterfactual generation can then be expressed as $\tilde{y}_{i}=g_{\tilde{y}_{<i}}\left(g_{y_{<i}}^{-1}(y_{i})\right)$ so that each counterfactual value can be calculated in sequence. If a label $y_{i}$ is missing, we use our trained predictor $q_{\phi}(y_{i}|x,y_{>i})$ to impute it, e.g Fig. 1 (right) for when $y_{E}$ is unobserved and we intervene on $y_{C}$.
@@ 4.1.1 Counterfactual Effectiveness
Table 1: Colour Morpho-MNIST: Log likelihoods of the child variables. Colour log likelihoods (, $\in(-\infty,\>2.239]$, intensity log likelihoods $\in(-\infty,\>-1.336]$.
Model Labelled MAE $\;p_{\theta}(f)\;$ $\;p_{\theta}(b)\;$ $\;p_{\theta}(i,t)\;$ $\;q(f|\tilde{x})\;$ $\;q(b|\tilde{x})\;$ $\;q(i,t|\tilde{x})\;$
@@ 4.1.2 Independence of Cause and Mechanism
Table 2: (left) Cause and mechanism experiment. (right) MIMIC-CXR interventions: rows are 10%, 20%, 30% labelled, all with counterfactual regularisation.
@@
$p_{\theta}(i,t)\;(\uparrow)$ $q(i,t|\tilde{x})\;(\uparrow)$ $||t-\tilde{t}||_{1}\;(\downarrow)$ $||i-\tilde{i}||_{1}\;(\downarrow)$
300 2700 -19.31 -29.40 0.094 0.320
600 2400 -16.77 -26.99 0.098 0.288
1200 1800 -18.80 -21.16 0.114 0.231
1500 1500 -15.28 -27.08 0.108 0.191
1800 1200 -21.05 -20.09 0.110 0.123
2400 600 -16.59 -19.97 0.117 0.106
2700 300 -14.72 -17.74 0.112 0.092
@@ 4.2 Counterfactuals for Medical Imaging Data
To evaluate our method on medical imaging data, we apply it to the MIMIC-CXR dataset (50k training, 20k test samples) [12, 9]. We assume the casual structure used in [7] with variables disease ($d$), age ($a$), sex ($s$), race ($r$), with the only non-independence being that $a$ causes $d$ (Fig. 2b). For disease, we use the presence of pleural effusion as a binary variable and we train models using 10%, 20%, 30%, 40% and 50% of the total labels for each of the three models (Supervised, SSL, Flexible). As the causal structure is simpler, we measure performance by intervening on each variable separately before measuring the accuracy via the ROCAUC for the discrete variables, and the MAE for age.
@@
From the cells on the diagonal of Table 4.1.2 (right), we see that our method tends to improve upon the supervised approach with regards to implementing interventions. The other cells are essentially a measure of reconstruction quality, since they involve evaluating variables that have not been intervened on. As such, the closeness of these values for the various models suggests that the achieved counterfactual generation gains are primarily due to differences in the causal inference component. This indicates that it would be fruitful to focus future efforts on improving this section of the model. This holds for both SSL and Flexible, demonstrating that practitioners implementing our approach need not prioritise achieving full labelling for any given sample over collecting as many labels as possible, bolstering the usability of the model. Fig. 4a demonstrates the increased interventional accuracy provided by counterfactual regularisation. Moreover, as shown in Figure 4b, our model is able to exhibit clear visual changes for the various counterfactuals, indicating the numerical results are not due to minute changes undetectable to the human eye [2]. To build upon this, one avenue of future research would be to use this approach to generate additional training data for underrepresented populations in medical datasets and evaluate how this aids downstream tasks.
@@ 5 Conclusion
6 Acknowledgements
Yasin Ibrahim is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1). The authors also acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (http://dx.doi.org/10.5281/zenodo.22558).

### v2

@@ Abstract
Developing models that are capable of answering questions of the form “How would $x$ change if $y$ had been $z$?” is fundamental to advancing medical image analysis. Training causal generative models that address such counterfactual questions, though, currently requires that all relevant variables have been observed and that the corresponding labels are available in the training data. However, clinical data may not have complete records for all patients and state of the art causal generative models are unable to take full advantage of this. We thus develop, for the first time, a semi-supervised deep causal generative model that exploits the causal relationships between variables to maximise the use of all available data. We explore this in the setting where each sample is either fully labelled or fully unlabelled, as well as the more clinically realistic case of having different labels missing for each sample. We leverage techniques from causal inference to infer missing values and subsequently generate realistic counterfactuals, even for samples with incomplete labels. Code is available at: https://github.com/yi249/ssl-causal
@@ 1 Introduction
The deployment of deep learning models to real-world applications faces a variety of challenges [16], with many arguing that this is due to lack of causal considerations [4, 18]. A growing research area is the generation of counterfactuals (CFs), the manifestation of a sample in an alternative world where an upstream variable has been changed [17, 21, 10]. Such techniques are particularly useful in medical image analysis, where models are often hampered by lack of diversity in training data [5], so methods to generate realistic synthetic samples from underrepresented classes are critical [13]. Incorporating structural causal equations into a deep learning framework has been shown to provide a powerful tool for counterfactual generation [17]. These ideas were extended by the development of a hierarchical VAE structure for greater image fidelity [7]. This method consists, however, of a separately trained generative model and structural causal model, represented by a directed acyclic graph (DAG), and hence, the two components are unable to leverage information from one another during training. Moreover, these methods rely on fully labelled samples, so are unable to use additional data where true values are unavailable for some (or all) variables of the causal graph.
@@ 2 Background
$\mathcal{M}=\langle V,U,\mathcal{F},P(U)\rangle$
where, $V=\{v_{1},\dots,v_{n}\}$ is the set of endogenous variables of interest, $U=\{u_{1},\dots,u_{n}\}$ is the set of exogenous (noise) variables, $P(U)$ is the prior distribution over them, and =$\mathcal{F}=\{f_{1},\dots,f_{n}\}$ is a set of functions assigning values to the endogenous variables. Moreover, we assume that each endogenous variable, $x_{i}$, is assigned deterministically by its direct causes, i.e. its parents $\text{pa}_{i}\subseteq V\setminus\{v_{i}\}$, and the corresponding noise variable, $u_{i}$ via the structural assignments,
@@
$P(V|\text{do}(v_{i}=a))=\prod_{j\neq i}p(v_{j}|pa_{j})\cdot\mathds{1}_{\{v_{i}=a\}}$
@@
    Abduction: Use the data to update the prior probability on the exogenous noise $p(U)$ to obtain $p(U|V)$
@@ 3 Methodology
An overview of our method is shown in Fig. 1. Herein, the endogenous variables consist of image, $x$, and variables, $y$; denoted by $y^{*}$ when observed and by $y^{\prime}$ when predicted. Latent variables $z=z_{1:L}$ make up part of the exogenous noise for $x$, modelled using a hierarchical latent structure, following [7, 20]. Our model extends this structure with a predictive part that infers the endogenous variables, $y$, enabling counterfactual generation in the case of missing labels. For clarity, we limit derivations to the case with a single cause variable, $y_{C}$, and effect variable, $y_{E}$, but this can be extended to any finite number of variables with any causal structure. The ELBO loss for fully labelled samples drawn from $\mathcal{D}_{L}$ is $\mathcal{S}(x,y)$:
@@
where $\mathcal{L}(x,y)$ is the ELBO for a conditional VAE and $p_{\theta}(\cdot)$ are (Gaussian) priors. For the unlabelled samples, drawn from $\mathcal{D}_{U}$, we minimise the loss:
@@
Under the current construction, the parent predictors, $q_{\phi}(y_{E}|x)$ and $q_{\phi}(y_{C}|x,y_{E})$, are only trained when the parent variables are unobserved. To ensure the model is able to learn a good representation of $y$, we include additional classification terms in the supervised loss [14], giving the total loss to train our model:
@@ 3.0.2 Counterfactual Generation
Once the generative model is trained, we use the predictive component $p_{\theta}(y_{i}|y_{<i})$ to generate counterfactuals, $\tilde{x}$. For abduction we require the structural assignments (1) to be invertible in $u$, so we encode $p_{\theta}(y_{i}|y_{<i})$ as an invertible $g_{y_{<i}}(u),\;u\sim p(u)$, parameterised by the causal parents $y_{<i}$ [17]. Counterfactual generation can then be expressed as $\tilde{y}_{i}=g_{\tilde{y}_{<i}}\left(g_{y_{<i}}^{-1}(y_{i})\right)$ so that each counterfactual value can be calculated in sequence. If a label $y_{i}$ is missing, we use our trained predictor $q_{\phi}(y_{i}|x,y_{>i})$ to impute it, e.g Fig. 1 (right) for when $y_{E}$ is unobserved and we intervene on $y_{C}$.
@@ 4.1.1 Counterfactual Effectiveness
Table 1: Colour Morpho-MNIST: Log likelihoods ()($\uparrow$)( ↑ ) of the child variables. Colour log likelihoods (, $\in(-\infty,\>2.239]$, intensity log likelihoods $\in(-\infty,\>-1.336]$.
Model Labelled $\;\text{MAE}\;(\downarrow)$ $\;p_{\theta}(f)\;$ $\;p_{\theta}(b)\;$ $\;p_{\theta}(i,t)\;$ $\;q(f|\tilde{x})\;$ $\;q(b|\tilde{x})\;$ $\;q(i,t|\tilde{x})\;$
@@ 4.1.2 Independence of Cause and Mechanism
Table 2: (left) Colour Morpho-MNIST: Cause and mechanism experiment. (right) MIMIC-CXR: For each intervention $do(\cdot)$, the 3 rows correspond to training with 10%, 20%, 30% of variables labelled, all using CF regularisation. Semi-supervision in both settings (SSL, Flexible) outperforms pure supervision (Sup.).
@@
$p_{\theta}(i,\!t)\;(\uparrow)$ $q(i,\!t|\tilde{x})\;(\uparrow)$ $|t\!-\!\tilde{t}|\;(\downarrow)$ $|i\!-\!\tilde{i}|\;(\downarrow)$
300 2700 -21.84 -29.47 0.089 0.293
600 2400 -18.89 -26.17 0.102 0.203
1200 1800 -18.01 -24.78 0.099 0.221
1500 1500 -19.40 -25.43 0.132 0.172
1800 1200 -16.27 -22.13 0.120 0.161
2400 600 -14.62 -19.32 0.146 0.093
2700 300 -14.15 -17.56 0.152 0.088
@@ 4.2 Counterfactuals for Medical Imaging Data
To evaluate our method on medical imaging data, we apply it to the MIMIC-CXR dataset (50k training, 20k test samples) [12, 9]. We assume the casual structure used in [7] with variables disease ($d$), age ($a$), sex ($s$), race ($r$), with the only non-independence being that $a$ causes $d$ (Fig. 2b). For disease, we use the presence of pleural effusion as a binary variable and we train models using 10%, 20%, 30%, 40% and 50% of the total labels for each of the three models (Supervised, SSL, Flexible). As the causal structure is simpler, we measure performance, over 3 seeds, by intervening on each variable separately before measuring the ROCAUC for the discrete variables and the MAE for age.
@@
From the cells on the diagonal of Table 4.1.2 (right), we see that our method tends to improve upon the supervised approach with regards to implementing interventions. The other cells are essentially a measure of reconstruction quality, since they involve evaluating variables that have not been intervened on. As such, the closeness of these values for the various models suggests that the achieved counterfactual generation gains are primarily due to differences in the causal inference component. This indicates that it would be fruitful to focus future efforts on improving this section of the model. This holds for both SSL and Flexible, demonstrating that practitioners implementing our approach need not prioritise achieving full labelling for any given sample over collecting as many labels as possible, bolstering the usability of the model. Fig. 4a demonstrates the increased interventional accuracy provided by CF regularisation. Moreover, as shown in Figure 4b, our model is able to exhibit clear visual changes for the various CFs, indicating the numerical results are not due to minute changes undetectable to the human eye [2]. To build upon this, an avenue of future research would be to use this approach to generate additional training data for underrepresented populations in medical datasets and evaluate how this aids downstream tasks.
@@ 5 Conclusion
{credits}
5.0.1 Acknowledgements
Yasin Ibrahim and Hermione Warr are supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1). The authors also acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (http://dx.doi.org/10.5281/zenodo.22558).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.